OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. -√2,3√2

OpenStudy (chaise):

If the solutions are: -√2 or 3√2 (x+√2)(x-3√2)=0 Foil it out: \[x^2-2\sqrt{2}x-6=0\]